OpenStudy (anonymous):

The volume of a rectangular box is found by multiplying its length, width, and height: V = lwh. A certain box has a volume of b3 + 3b2 - 4b - 12. Factor the four-term polynomial to find the dimensions of the box. Hint: Use factoring by grouping to factor the four term polynomial first. b - 3; b - 2; b - 2 b + 3; b - 2; b - 2 b + 3; b + 2; b - 2

OpenStudy (anonymous):

will u help me @Cass06

pooja195 (pooja195):

x-8=48^1/2 x=(48^1/2)+8 or x=square root of 48 +8

OpenStudy (anonymous):

whts the answer

pooja195 (pooja195):

it can be either they are both the same soo choose

OpenStudy (anonymous):

I say the 2nd one am I right

pooja195 (pooja195):

its much more easier to write it that way so yes

OpenStudy (anonymous):

okay thank u

pooja195 (pooja195):

no problem